# Learning Encodings by Maximizing State Distinguishability: Variational Quantum Error Correction

### v1

@@ Appendix D Implementation Details and Setup
In this appendix, we summarize details regarding the actual experimental realization and implementation. We introduce a new type of ansatz we used for both the encoding and recovery operations in Section D.1. In Section D.2 we highlight some implementation details and also discuss some algorithmic hyperparameters.
@@ D.1 Randomized Entangling Ansätze
Figure 13: An instance of the randomized entangling ansatz (RAG) on $n$ qubits. It consists of an initial block of parameterized single-qubit gates $V$ and $d$ blocks of successive entangling operations. The entangling blocks themselves consist of an optionally parametrized two-qubit gate $W$, and parameterized single-qubit operations $V$ on the two involved qubits. The control and target for each two-qubit block are determined uniformly at random, i.e. the exact instance of an ansatz is defined by the random seed that was used for generating it.
In Fig. 13 we show an instance of such an RAG ansatz for the encoding operation. We note that the same ansatz can also be used for the recovery operations, potentially incorporating the additional recovery register. The general outline is defined by a parameterized single-qubit operation $V$ and an optionally also parameterized two-qubit operation $W$. Throughout this paper, we default to
@@ D.2 Hyperparameters and Python Module
This is the central component of the VarQEC pipeline and integrates all other components. It imports the respective models, circuit ansätze, error channels, and loss functions. It currently makes use of the L-BFGS optimizer [58] with a history size of $100$ and $10$ iterations per training epoch, but can easily be adapted to support other optimizer instances. The training script allows for the isolated training of encoding operations with the distinguishability loss, but also supports the successive training of recovery operations with the fidelity loss. The testing script allows to perform a detailed analysis of the trained models, but also implements several known error correction codes as a baseline. Furthermore, it also allows to evaluate the potential code distance introduced in Section B.3.
@@
These are one of the main contributions of this paper, and are therefore implemented in a modular function, to enable usage also in other setups. For the fidelity loss, we only provide a simplified implementation based on the QVECTOR approach, as described in Appendix C. For the distinguishability loss, we provide two separately finetuned versions, one to reduce memory overhead on the expense of additional time, and one that is adapted the other way around.
@@
We provide two different methods to prepare the initial states, corresponding to the exact and approximate formulation of the respective loss functions. The first one initializes an adjustable number of Haar-random states. This is done by sampling the respective number of unitaries of size $2^{k}\times 2^{k}$ randomly from the unitary group $\mathcal{U}(2^{k})$ [68], which conceptually is supported for an arbitrary number of data qubits $k$. The second method prepares states from a two-design. To enhance the experimental efficiency, which is crucial as the loss function is trained on the two-design approximations, we implement the respective state initializations in a parameterized way to allow for batch processing. One single qubit, i.e. $k=1$, we place an arbitrary single-qubit rotation $U_{3}$, which is initialized with the values denoted in Table 3 produces the desired states. We implement the same methodology also for $k=2$ qubits, but we sketched in Section B.4 how this also could be extended to larger systems.
|$\ket{0}$ |$\ket{1}$ $\ket{+}$ $\ket{-}$ |$\ket{+i}$ |$\ket{-i}$
$U_{3}(0.0,0.0,0.0)$ $\leavevmode\nobreak\ U_{3}(\pi,0.0,\pi)\leavevmode\nobreak$ $\leavevmode\nobreak\ U_{3}(\frac{\pi}{2},0.0,\pi)\leavevmode\nobreak$ $U_{3}(\frac{\pi}{2},-\pi,-\pi)$ $\leavevmode\nobreak\ U_{3}(\frac{\pi}{2},\frac{\pi}{2},-\pi)\leavevmode\nobreak$ $U_{3}(\frac{\pi}{2},-\frac{\pi}{2},-\pi)$
@@
The circuit ansatz determines the layout and allocation of trainable parameters that are trained to be the encoding or, respectively, recovery operation. In the VarQEC module, we implement the randomized entangling ansatz (REA) detailed in Section D.1, but it is also readily extensible to other custom ansätze. The module enables setting a restricted connectivity as in Section E.4, which selects target and control pairs only for qubits that are physically connected.
@@
The noise channels are implemented and simulated using the default.mixed simulator of PennyLane. This includes all the noise channels that were detailed in Section A.2. It is also straightforward to extend the module with custom noise channels, or directly load hardware properties like $T_{1}$ and $T_{2}$ times.
@@ Appendix E Extended Empirical Results
In this appendix, we provide additional empirical evaluations to further substantiate the findings presented in Section 5. We explore the performance of the proposed VarQEC method under various noise channels, assess its scalability to multiple logical qubits, further examine its stability, and analyze the resilience under restricted connectivity.
@@ E.1 Additional Noise Channels
(b) Thermal relaxation noise of duration $t=10.0$us with $T_{1}=200$us and $T_{2}=100$us.
@@
Figure 14: Distinguishability loss for different QEC codes on (a) bit-flip, (b) thermal relaxation, (c) amplitude damping, and (d) consecutive amplitude and phase damping noise. All bars visualize the average-case loss $\mathcal{D}$ and worst-case loss $\overline{\mathcal{D}}$. Additionally, the two-design approximations $\mathcal{D}_{\mathcal{S}},\overline{\mathcal{D}}_{\mathcal{S}}$ are visualized as dotted lines if the approximation was not exact. For the noise models, the unencoded baseline, the [[$5,1,3$]] perfect code [53], optionally the [[$3,1,1$]] bit-flip [85] and [[$4,1,2$]] approximate QEC code [55], and VarQEC codes with $n=3$ to $n=7$ physical qubit are compared. All VarQEC ansätze are trained for $10$ epochs and contain $(n-1)\cdot(n-2)$ two-qubit blocks. Results on symmetric and asymmetric depolarizing noise can be found in Fig. 5.
We first extend our analysis to learning encoding operations for additional noise models. The results for bit-flip, thermal relaxation, amplitude damping, and consecutive amplitude and phase damping channels are given in Fig. 14. For the bit-flip noise in Fig. 14(a) additional to the [[$5,1,3$]] perfect code we also consider the [[$3,1,1$]] bit-flip code [85]. Confirming our observations from the main part, the [[$3,1$]] VarQEC code performs on par with the bit-flip code, and the [[$5,1$]] VarQEC code performs on par with the perfect code. Moreover, increasing the number of physical qubits to $n=7$ allows for further reduction of the distinguishability loss. For the thermal relaxation channel in Fig. 14(b), results are similar to those in depolarizing noise in Fig. 5(a), but the VarQEC codes manage to have a small edge over the perfect code. A similar connection can be observed for the amplitude and phase damping noise in Fig. 14(d). for isolated amplitude damping noise in Fig. 14(c) we additionally compare the [[$4,1,2$]] approximate QEC code [55]. Also, here increasing the number of physical qubits in the VarQEC codes significantly reduces the achieved distinguishability loss further.
@@
(b) Thermal relaxation noise of duration $t=10.0$us with $T_{1}=200$us, $T_{2}=100$us.
@@
Figure 15: Fidelity loss for different QEC codes and respective recovery operations on (a) bit-flip, (b) thermal relaxation, (c) amplitude damping, and (d) consecutive amplitude and phase damping noise. All bars visualize the average-case loss $\mathcal{F}$ and worst-case loss $\overline{\mathcal{F}}$. For the noise models, the unencoded baseline, and the [[$5,1,3$]] perfect code is visualized, optionally extended by the [[$3,1,1$]] bit-flip and [[$4,1,2$]] approximate QEC code. Additionally, we compare three instances of VarQEC codes, where [[$5,1$]] employs a random encoding, [[$5,1$]]∗∗ employs a encoding that has converged to a locally optimal distinguishability loss, and the encoding of [[$5,1$]] has converged to a global optimum. The recovery operations are successively trained following Eq. 22. For the rightmost model, the encoding and recovery operations have been trained end-to-end with the QVECTOR approach [46] as described in Appendix C. The recovery ansätze are trained using the fidelity loss for $50$ epochs and contain $200$ two-qubit blocks. As we consider a single-recovery scenario, no separate recovery register was used, i.e. the recovery operation acts solely on the $n$-qubit encoded and disturbed state [46]. The distinguishability loss for all encodings is summarized in Table 4. Results on symmetric and asymmetric depolarizing noise can be found in Fig. 7.
Base- [[$5,1,3$]] [[$5,1$]] [[$5,1$]]∗∗ [[$5,1$]] [[$5,1$]]
@@
Table 4: Worst-case distinguishability loss $\overline{\mathcal{D}}$ exhibited by the encodings for the models in Fig. 15. For the baseline and [[$5,1,3$]] code, this is evaluated on the unencoded baseline and static encoding operation, respectively. For the VarQEC codes, this corresponds to an untrained encoding for [[$5,1$]], to an local optimum for [[$5,1$]]∗∗, and to a global optimum for [[$5,1$]]. For the [[$5,1$]] QVECTOR code, the same ansatz with $12$ two-qubit blocks as for the VarQEC codes is employed, but training was only implicitly conducted via the fidelity loss. Additionally, we again report the worst-case fidelity loss $\overline{\mathcal{F}}$ (third line) and the respective upper (second) and upper bounds (fourth line) following Theorems 2 and 1 for the VarQEC models.
Building upon this, we train respective recovery operations and evaluate the fidelity losses in Fig. 7. The distinguishability losses of the underlying encoding circuits and the implied bounds are summarized in Table 4. In agreement with the results from the main part, we can observe that a lower distinguishability loss allows for a higher-fidelity recovery operation. On all considered instances, we can observe that the optimal VarQEC codes outperform the unencoded baseline, the [[$5,1,3$]] perfect code, and also the [[$5,1$]] code trained with the QVECTOR approach. In Section B.3 we determine the [[$5,1$]] VarQEC code trained on amplitude damping noise to exhibit a potential approximate code distance of $d=2$. This aligns with the code distance of the [[$4,1,2$]] approximate QEC code, which is known to be especially suited for the amplitude damping channel [55]. We again want to emphasize that the QVECTOR approach employs the same circuit ansätze as the VarQEC codes. For bit-flip and amplitude damping noise, the [[$5,1$]] QVECTOR codes also significantly improve upon the unencoded baseline performance, which is in agreement with the results from Ref. [46]. However, already on these and especially on the more complex thermal relaxation noise channel, this end-to-end ansatz produces results that are inferior to the VarQEC procedure.
@@ E.2 Multiple Data Qubits
We now assess the extension of the previous experiments to the setup with multi-qubit logical states, in particular $k=2$. For this two-qubit setup, we enumerate elements of a two-design in Section B.4. In Fig. 16, we report the distinguishability loss for a variety of such setups on both symmetric and asymmetric depolarizing noise. We include the unencoded baseline, which in this context corresponds to a two-qubit state. We additionally compare to the [[$4,2,2$]] QEC code [97] and the [[$8,2,3$]] QEC code [71]. Using our VarQEC routine, we train codes with sizes from $n=4$ to $n=8$ physical qubits.
@@
(b) Asymmetric depolarizing noise with noise strength $p=0.1$ and asymmetry $c=0.5$ following Eq. 41.
Figure 16: Distinguishability loss for different QEC codes with two logical qubits, i.e. $k=2$, on (a) depolarizing and (b) asymmetric depolarizing noise. All bars visualize the average-case loss $\mathcal{D}$ and worst-case loss $\overline{\mathcal{D}}$. Additionally, the two-design approximations $\mathcal{D}_{\mathcal{S}},\overline{\mathcal{D}}_{\mathcal{S}}$ are visualized as dotted lines if the approximation was not exact. For both noise models, the unencoded baseline, the [[$4,2,2$]] QEC code [97], the [[$8,2,3$]] QEC code [71], and VarQEC codes with $n=4$, $n=6$, and $n=8$ physical qubits are compared. All VarQEC ansätze are trained for $10$ epochs and contain $n\cdot(n-1)$ two-qubit blocks. Results on only a single data qubit can be found in Fig. 5.
The results in on the symmetric depolarizing channel in Fig. 16(a) indicate that for this setup, the [[$8,2$]] VarQEC are not fully competitive with the [[$8,2,3$]] QEC code, which, similar to the perfect code on a single logical qubit, is the smallest code with distance $d=3$ on two logical qubits. However, we assume this gap could be eliminated by resorting to a more sophisticated ansatz construction. On the contrary, the results on asymmetric depolarizing noise in Fig. 16(b) demonstrate that the VarQEC codes reduce the distinguishability loss compared to the standard QEC codes.
Overall, we want to recall that in QEC it is usually not pursued to encode a high number of logical qubits into one code patch. Rather, one typically aims at encoding a small number of physical qubits per code, and then establishing interactions between these patches via some fault-tolerant protocols. The prime example would be lattice surgery for surface codes, where only a single logical qubit is encoded per patch [43]. The development of such protocols and operations for the VarQEC procedure is left for future work.

### v2

@@ Appendix D Implementation Details and Setup
In this appendix, we summarize details regarding the actual experimental realization and implementation. We introduce a new type of ansatz we used for both the encoding and recovery operations in section˜D.1. In section˜D.2 we highlight some implementation details and also discuss some algorithmic hyperparameters.
@@ D.1 Randomized Entangling Ansätze
Figure 13: An instance of the REA on $n$ qubits. It consists of an initial block of parameterized single-qubit gates $V$ and $d$ blocks of successive entangling operations. The entangling blocks themselves consist of an optionally parametrized two-qubit gate $W$, and parameterized single-qubit operations $V$ on the two involved qubits. The control and target for each two-qubit block are determined uniformly at random, i.e. the exact instance of an ansatz is defined by the random seed that was used for generating it.
In figure˜13 we show an instance of such a REA ansatz for the encoding operation. We note that the same ansatz can also be used for the recovery operations, potentially incorporating the additional recovery register. The general outline is defined by a parameterized single-qubit operation $V$ and an optionally also parameterized two-qubit operation $W$. Throughout this paper, we default to
@@ D.2 Hyperparameters and Python Module
This is the central component of the VarQEC pipeline and integrates all other components. It imports the respective models, circuit ansätze, error channels, and loss functions. It currently makes use of the L-BFGS optimizer [63] with a history size of $100$ and $10$ iterations per training epoch, but can easily be adapted to support other optimizer instances. The training script allows for the isolated training of encoding operations with the distinguishability loss, but also supports the successive training of recovery operations with the fidelity loss. The testing script allows to perform a detailed analysis of the trained models, but also implements several known error correction codes as a baseline. Furthermore, it also allows to evaluate the potential code distance introduced in section˜B.3.
@@
These are one of the main contributions of this paper, and are therefore implemented in a modular function, to enable usage also in other setups. For the fidelity loss, we only provide a simplified implementation based on the QVECTOR approach, as described in Appendix˜C. For the distinguishability loss, we provide two separately finetuned versions, one to reduce memory overhead on the expense of additional time, and one that is adapted the other way around.
@@
We provide two different methods to prepare the initial states, corresponding to the exact and approximate formulation of the respective loss functions. The first one initializes an adjustable number of Haar-random states. This is done by sampling the respective number of unitaries of size $2^{k}\times 2^{k}$ randomly from the unitary group $\mathcal{U}(2^{k})$ [74], which conceptually is supported for an arbitrary number of data qubits $k$. The second method prepares states from a two-design. To enhance the experimental efficiency, which is crucial as the loss function is trained on the two-design approximations, we implement the respective state initializations in a parameterized way to allow for batch processing. One single qubit, i.e. $k=1$, we place an arbitrary single-qubit rotation $U_{3}$, which is initialized with the values denoted in table˜3 produces the desired states. We implement the same methodology also for $k=2$ qubits, but we sketched in section˜B.4 how this also could be extended to larger systems.
[…]
$U_{3}(0.0,0.0,0.0)$ $\penalty 10000\ U_{3}(\pi,0.0,\pi)\penalty 10000\$$\penalty 10000\ U_{3}(\frac{\pi}{2},0.0,\pi)\penalty 10000\$$U_{3}(\frac{\pi}{2},-\pi,-\pi)$ $\penalty 10000\ U_{3}(\frac{\pi}{2},\frac{\pi}{2},-\pi)\penalty 10000\$$U_{3}(\frac{\pi}{2},-\frac{\pi}{2},-\pi)$
@@
The circuit ansatz determines the layout and allocation of trainable parameters that are trained to be the encoding or, respectively, recovery operation. In the VarQEC module, we implement the REA detailed in section˜D.1, but it is also readily extensible to other custom ansätze. The module enables setting a restricted connectivity as in section˜E.4, which selects target and control pairs only for qubits that are physically connected.
@@
The noise channels are implemented and simulated using the default.mixed simulator of PennyLane. This includes all the noise channels that were detailed in section˜A.2. It is also straightforward to extend the module with custom noise channels, or directly load hardware properties like $T_{1}$ and $T_{2}$ times.
@@ Appendix E Extended Empirical Results
In this appendix, we provide additional empirical evaluations to further substantiate the findings presented in section˜5. We explore the performance of the proposed VarQEC method under various noise channels, assess its scalability to multiple logical qubits, further examine its stability, and analyze the resilience under restricted connectivity.
@@ E.1 Additional Noise Channels
(b) Thermal relaxation noise of duration t=10.0t=10.0us with $T_{1}=200$us and $T_{2}=100$us.
@@
Figure 14: Distinguishability loss for different QEC codes on (a) bit-flip, (b) thermal relaxation, (c) amplitude damping, and (d) consecutive amplitude and phase damping noise. All bars visualize the average-case loss $\mathcal{D}$ and worst-case loss $\overline{\mathcal{D}}$. Additionally, the two-design approximations $\mathcal{D}_{\mathcal{S}},\overline{\mathcal{D}}_{\mathcal{S}}$ are visualized as dotted lines if the approximation was not exact. For the noise models, the unencoded baseline, the $[[5,1,3]]$ perfect code [57], optionally the $[[3,1,1]]$ bit-flip [92] and $[[4,1,2]]$ approximate QEC code [59], and VarQEC codes with $n=3$ to $n=7$ physical qubit are compared. All VarQEC ansätze are trained for $10$ epochs and contain $(n-1)\cdot(n-2)$ two-qubit blocks. Results on symmetric and asymmetric depolarizing noise can be found in figure˜5.
We first extend our analysis to learning encoding operations for additional noise models. The results for bit-flip, thermal relaxation, amplitude damping, and consecutive amplitude and phase damping channels are given in figure˜14. For the bit-flip noise in figure˜14(a) additional to the $[[5,1,3]]$ perfect code we also consider the $[[3,1,1]]$ bit-flip code [92]. Confirming our observations from the main part, the $((3,2))$ VarQEC code performs on par with the bit-flip code, and the $((5,2))$ VarQEC code performs on par with the perfect code. Moreover, increasing the number of physical qubits to $n=7$ allows for further reduction of the distinguishability loss. For the thermal relaxation channel in figure˜14(b), results are similar to those in depolarizing noise in figure˜5(a), but the VarQEC codes manage to have a small edge over the perfect code. A similar connection can be observed for the amplitude and phase damping noise in figure˜14(d). for isolated amplitude damping noise in figure˜14(c) we additionally compare the $[[4,1,2]]$ approximate QEC code [59]. Also, here increasing the number of physical qubits in the VarQEC codes significantly reduces the achieved distinguishability loss further.
@@
(b) Thermal relaxation noise of duration t=10.0t=10.0us with $T_{1}=200$us, $T_{2}=100$us.
@@
Figure 15: Fidelity loss for different QEC codes and respective recovery operations on (a) bit-flip, (b) thermal relaxation, (c) amplitude damping, and (d) consecutive amplitude and phase damping noise. All bars visualize the average-case loss $\mathcal{F}$ and worst-case loss $\overline{\mathcal{F}}$. For the noise models, the unencoded baseline, and the $[[5,1,3]]$ perfect code is visualized, optionally extended by the $[[3,1,1]]$ bit-flip and $[[4,1,2]]$ approximate QEC code. Additionally, we compare three instances of VarQEC codes, where $((5,2))^{*}$ employs a random encoding, $((5,2))^{**}$ employs a encoding that has converged to a locally optimal distinguishability loss, and the encoding of $((5,2))$ has converged to a global optimum. The recovery operations are successively trained following equation 26. For the rightmost model, the encoding and recovery operations have been trained end-to-end with the QVECTOR approach [50] as described in Appendix˜C. The recovery ansätze are trained using the fidelity loss for $50$ epochs and contain $200$ two-qubit blocks. As we consider a single-recovery scenario, no separate recovery register was used, i.e. the recovery operation acts solely on the $n$-qubit encoded and disturbed state [50]. The distinguishability loss for all encodings is summarized in table˜4. Results on symmetric and asymmetric depolarizing noise can be found in figure˜7.
Base- $[[5,1,3]]$ $^{*} ((5,2))((5,2))^{**}$ $((5,2))$ $((5,2))$
@@
Table 4: Worst-case distinguishability loss $\overline{\mathcal{D}}$ exhibited by the encodings for the models in figure˜15. For the baseline and $[[5,1,3]]$ code, this is evaluated on the unencoded baseline and static encoding operation, respectively. For the VarQEC codes, this corresponds to an untrained encoding for $((5,2))^{*}$, to an local optimum for $((5,2))^{**}$, and to a global optimum for $((5,2))$. For the $((5,2))$ QVECTOR code, the same ansatz with $12$ two-qubit blocks as for the VarQEC codes is employed, but training was only implicitly conducted via the fidelity loss. Additionally, we again report the worst-case fidelity loss $\overline{\mathcal{F}}$ (third line) and the respective upper (second) and upper bounds (fourth line) following Theorems˜2 and 1 for the VarQEC models.
(a) Depolarizing noise with noise strength $p=0.1$.
(b) Asymmetric depolarizing noise with noise strength $p=0.1$ and asymmetry $c=0.5$ following equation˜45.
Figure 16: Distinguishability loss for different QEC codes with two logical qubits, i.e. $k=2$, on (a) depolarizing and (b) asymmetric depolarizing noise. All bars visualize the average-case loss $\mathcal{D}$ and worst-case loss $\overline{\mathcal{D}}$. Additionally, the two-design approximations $\mathcal{D}_{\mathcal{S}},\overline{\mathcal{D}}_{\mathcal{S}}$ are visualized as dotted lines if the approximation was not exact. For both noise models, the unencoded baseline, the $[[4,2,2]]$ QEC code [105], the $[[8,2,3]]$ QEC code [78], and VarQEC codes with $n=4$, $n=6$, and $n=8$ physical qubits are compared. All VarQEC ansätze are trained for $10$ epochs and contain $n\cdot(n-1)$ two-qubit blocks. Results on only a single data qubit can be found in figure˜5.
Building upon this, we train respective recovery operations and evaluate the fidelity losses in figure˜7. The distinguishability losses of the underlying encoding circuits and the implied bounds are summarized in table˜4. In agreement with the results from the main part, we can observe that a lower distinguishability loss allows for a higher-fidelity recovery operation. On all considered instances, we can observe that the optimal VarQEC codes outperform the unencoded baseline, the $[[5,1,3]]$ perfect code, and also the $((5,2))$ code trained with the QVECTOR approach. In section˜B.3 we determine the $((5,2))$ VarQEC code trained on amplitude damping noise to exhibit a potential approximate code distance of $d=2$. This aligns with the code distance of the $[[4,1,2]]$ approximate QEC code, which is known to be especially suited for the amplitude damping channel [59]. We again want to emphasize that the QVECTOR approach employs the same circuit ansätze as the VarQEC codes. For bit-flip and amplitude damping noise, the $((5,2))$ QVECTOR codes also significantly improve upon the unencoded baseline performance, which is in agreement with the results from Ref. [50]. However, already on these and especially on the more complex thermal relaxation noise channel, this end-to-end ansatz produces results that are inferior to the VarQEC procedure.
@@ E.2 Multiple Data Qubits
We now assess the extension of the previous experiments to the setup with multi-qubit logical states, in particular $k=2$. For this two-qubit setup, we enumerate elements of a two-design in section˜B.4. In figure˜16, we report the distinguishability loss for a variety of such setups on both symmetric and asymmetric depolarizing noise. We include the unencoded baseline, which in this context corresponds to a two-qubit state. We additionally compare to the $[[4,2,2]]$ QEC code [105] and the $[[8,2,3]]$ QEC code [78]. Using our VarQEC routine, we train codes with sizes from $n=4$ to $n=8$ physical qubits.
The results in on the symmetric depolarizing channel in figure˜16(a) indicate that for this setup, the $((8,4))$ VarQEC are not fully competitive with the $[[8,2,3]]$ QEC code, which, similar to the perfect code on a single logical qubit, is the smallest code with distance $d=3$ on two logical qubits. However, we assume this gap could be eliminated by resorting to a more sophisticated ansatz construction. On the contrary, the results on asymmetric depolarizing noise in figure˜16(b) demonstrate that the VarQEC codes reduce the distinguishability loss compared to the standard QEC codes.
We recall that in QEC it is standard to encode only a few logical qubits in a single code patch. Surface codes typically store one logical qubit per patch [35], while modern Q-LDPC codes can encode around a dozen logical qubits [13]. Fault-tolerant interactions between such patches are implemented using protocols like lattice surgery [47]. Developing analogous protocols and operations for the VarQEC procedure is left for future work.
@@
(b) Asymmetric depolarizing noise with noise strength $p=0.1$ and asymmetry $c=0.5$ following equation˜45.
Figure 17: Stability of standard QEC and VarQEC codes with increasing noise strength for (a) symmetric and (b) asymmetric depolarizing noise. All data points represent the respective worst-case distinguishability loss $\overline{\mathcal{D}}$. The $((5,2))^{*}$ VarQEC code is always the same static one trained on a noise strength $p=0.1$. The other $((5,2))$ VarQEC code is dynamically trained on the respective denoted noise configuration.
(a) Depolarizing noise with noise strength $p=0.1$ on n1n-1 of the qubits, and with a noise strength as denoted on the $x$-axis on the remaining qubit. For the QEC codes, the deviating qubit is associated with the original data qubit, while the unencoded baseline employs the best available wire.
Figure 18: Stability of standard QEC and VarQEC codes in setups of varying (a) noise strength on different qubits and (b) noise channels. All data points represent the respective worst-case distinguishability loss $\overline{\mathcal{D}}$ and optionally average-case loss $\mathcal{D}$. The $((5,2))^{*}$ VarQEC code is always the same static one trained on depolarizing noise with noise strength $p=0.1$ on all qubits. The other $((5,2))$ VarQEC code is dynamically trained on the respective noise configuration.